Write an equation containing the given points and parallel to the line
(-7,4); 7x=3y+5

Found 2 solutions by Mathtut, solver91311:
Answer by Mathtut(3670)

any line parallel to 7x=3y+5 will have the same slope.....so the first step is to find the slope of this line 3y=7x-5---->y=(7/3)x-5/3.......so the slope is 7/3
:
Now using the point slope formula with (-7,4) as our point
:
y-4=(7/3)(x+7)
:
y-4=7/3x+49/3
:
3y-12=7x+49.........multiplied all terms by 3
:
7x+3y=61

Answer by solver91311(24713)



Put your equation into slope-intercept form,

Determine the slope of your given line by inspection of the coefficient on x.

The slope of any line parallel to the given line must have the same slope because:



So use the point-slope form of the equation of a line, the slope of the given line, and the coordinates of the given point to derive the desired equation:



After substituting the values, I recommend that you put the equation into standard form so that it is like the given equation.

John